# CoSteer: Collaborative Decoding-Time Personalization via Local Delta Steering

### v1

@@ 1 Introduction
The rapid development of large language models (LLMs) has significantly enhanced natural language processing, enabling these models to understand context and generate coherent text effectively [1, 2]. This progress has sparked interest in personalization, where AI systems move beyond generic content to tailor interactions based on individual user profiles. Personalized generation refers to creating content customized through analysis of user-specific attributes like linguistic patterns, interaction histories, and contextual preferences [3]. This approach enables user-specific outputs that maintain contextual relevance, with applications in personalized recommender systems [4, 5, 6], adaptive dialogue agents [7], and customized content creation platforms [8].
Existing personalized generation approaches primarily fall into two paradigms. The first involves training-based methods, which leverage user data to tailor models using two main strategies. Parameter-efficient architectures, such as PLoRA [9], employ a shared LoRA module across users to strike a balance between personalization and resource efficiency. In contrast, individualized adapters, as seen in UserAdapter [10], optimize for per-user customization, albeit at a higher computational cost. Additionally, multi-objective reinforcement learning frameworks [11, 12] redefine personalization as the alignment of generic model capabilities with user-specific patterns, achieved through reward shaping. The second paradigm comprises tuning-free methods, which avoid model updates by adapting to user context during inference. This approach is further divided into two techniques. Prompt engineering implicitly encodes user profiles through curated prompts. For example, Cue-CoT [13] integrates user-specific reasoning into system prompts, while PAG [14] retrieves historical interactions to construct context-aware prompts. On the other hand, inference-time optimization techniques, exemplified by PAD [15] and Drift [16], actively adjust token generation probabilities by manipulating logit differences. Unlike prompt-based methods that inject semantic context, these techniques constrain output distributions to better align with user preferences.
@@
  • Extensive Experimental Validation: We provide substantial experimental evidence demonstrating the effectiveness of our proposed method. This finding is expected to illuminate and guide future research efforts.
@@ 2 Related work
Current mainstream approaches to personalization predominantly focus on training-time adaptation, leveraging user-specific data to tailor models through parameter-efficient fine-tuning (PEFT) or reinforcement learning-based alignment. A subset of work explores one-PEFT-for-all-users strategies [9, 17, 18, 19], where a shared set of lightweight parameters (e.g., adapters or conditional batch normalization modules) is optimized to generalize across diverse user preferences. Conversely, the one-PEFT-per-user paradigm prioritizes individualized tuning through isolated parameter sets, enhancing personalization while preserving privacy by avoiding cross-user data leakage [10, 20, 21, 22]. Beyond PEFT, RL-based approaches have gained traction for aligning models with user preferences through reward-driven optimization. Recent work has further formalized personalization as a multi-objective reinforcement learning (MORL) problem: methods like MORLHF [12] and MODPO [11] train distinct reward models for different objectives and merge them during optimization, while alternative frameworks such as Personalized Soups [23], Reward Soups [24], MOD [25], and PAD [15] dynamically combine policies from multiple trained models during the decoding phase. This series of work requires significant computational resources and relies on static datasets, which cannot meet the real-time personalization needs of users.
@@
Methods like Linear Alignment [26]and Contrastive Decoding [27] pioneer logit steering mechanisms that dynamically adjust token distributions during generation. By computing differential signals between personalized and generic outputs on edge devices, these approaches eliminate dependency on retraining while preserving privacy through localized computation. CoS [28] follows this approach by using a tunable parameter $\lambda$ to scale the impact of personal information, thereby achieving controllable personalization. Amulet [29] uses online learning algorithms to iteratively optimize the logits distribution, precisely aligning with the user’s preferences. However, this series of methods requires explicitly transmitting user information to the LLM, which poses a risk of privacy leakage and hinders the practical deployment of these methods.
@@
The computational constraints of edge devices have driven innovative paradigms in collaborative generation between cloud-based LLMs and local small language models. Existing collaborative generation approaches predominantly focus on two technical routes: (1) assistive alignment, where local SLMs are trained to augment the LLM’s user preference alignment during inference, and (2) reward-guided decoding, where lightweight reward models provide preference signals through reward-guided decoding. For instance, Aligner [30] leverages natural language feedback generated by the SLM to directly inject user-specific linguistic patterns into the LLM’s decoding process, while Expo [31] achieves preference alignment through linear interpolation of layer-wise weights between LLMs and SLMs. Recent advancements further demonstrate that training specialized small reward models can effectively steer LLM outputs toward personalized objectives [32, 33]. Among these, Proxy-tuning [34] and Cogensis [35] are closest to our design. Proxy-tuning modifies the LLM’s logits distribution by contrasting pre-trained and fine-tuned SLM outputs, whereas Cogensis employs a learned fusion network to combine logits from both models. In contrast to these existing methods, our proposed CoSteer framework establishes a completely tuning-free collaboration mechanism, eliminating the need for training.
@@ 3.1.2 Online learning
Inspired by the recent work Amulet [29], our study employs the widely studied Follow-The-Regularized-Leader (FTRL) [36] as the online learning algorithm. The FTRL is the FTL algorithm with a regularizer term added, which significantly improves both algorithmic stability and convergence properties. Typically, the policy optimization process of FTRL at iteration $t$ can be mathematically represented as:
@@ 3.2 Decoding-time personalization via local delta steering
Series of recent works, including Contrastive Decoding [27] and Linear Alignment [26], have demonstrated that the difference in logits before and after incorporating specific context into the model can serve as a direction for optimization [28, 29]. Thus from the online learning perspective, their utility function can be expressed as follows:
@@
where $s$ denotes the current token sequence, $P_{LLM}$/$\pi$ denotes the LLM’s distribution, and $P_{SLM}$/$\pi^{*}$ the SLM’s distribution. Hyperparameters $\alpha,\beta$ adjust the influence of each optimization component.
@@
By substituting Equation 4 into 5, and then 5 into 6, we can iteratively optimize the strategy. However, the current iterative computation incurs significant overhead, which is not feasible for our personalization scenarios that require rapid response. Therefore, we deduce the closed-form solution of Equation 6, as shown in Equation 7, to significantly reduce the additional computational cost. The derivation of the closed-form solution can be found in Appendix A.

### v2

@@ 1 Introduction
The rapid development of large language models (LLMs) has significantly enhanced natural language processing, enabling these models to understand context and generate coherent text effectively (Zhao et al., 2023; Gu et al., 2025; Huang et al., 2025). This progress has sparked interest in personalization, where AI systems move beyond generic content to tailor interactions based on individual user profiles. Personalized generation refers to creating content customized through analysis of user-specific attributes like linguistic patterns, interaction histories, and contextual preferences (Xu et al., 2025; Liu et al., 2025a). This approach enables user-specific outputs that maintain contextual relevance, with applications in personalized recommender systems (Lyu et al., 2023; Zhang et al., 2024a; Yu et al., 2025), adaptive dialogue agents (Wu et al., 2025), and customized content creation platforms (Mysore et al., 2024).
Existing personalized generation approaches primarily fall into two paradigms. The first involves training-based methods, which leverage user data to tailor models using two main strategies. Parameter-efficient architectures, such as PLoRA (Zhang et al., 2025a), employ a shared LoRA module across users to strike a balance between personalization and resource efficiency. In contrast, individualized adapters, as seen in UserAdapter (Zhong et al., 2021), optimize for per-user customization, albeit at a higher computational cost. Additionally, multi-objective reinforcement learning frameworks (Zhou et al., 2024; Wu et al., 2023) redefine personalization as the alignment of generic model capabilities with user-specific patterns, achieved through reward shaping. The second paradigm comprises tuning-free methods, which avoid model updates by adapting to user context during inference. This approach is further divided into two techniques. Prompt engineering implicitly encodes user profiles through curated prompts. For example, Cue-CoT (Wang et al., 2023a) integrates user-specific reasoning into system prompts, while PAG (Richardson et al., 2023) retrieves historical interactions to construct context-aware prompts. On the other hand, inference-time optimization techniques, exemplified by PAD (Chen et al., 2025) and Drift (Kim et al., 2025), actively adjust token generation probabilities by manipulating logit differences. Unlike prompt-based methods that inject semantic context, these techniques constrain output distributions to better align with user preferences.
@@
  • Comprehensive and Realistic Validation: We provide substantial experimental evidence demonstrating the effectiveness of our proposed method. Our evaluation further extends to address critical real-world challenges, confirming the framework’s practical value.
@@ 2 Related work
Current mainstream approaches to personalization predominantly focus on training-time adaptation, leveraging user-specific data to tailor models through parameter-efficient fine-tuning (PEFT) or reinforcement learning-based alignment. A subset of work explores one-PEFT-for-all-users strategies (Zhang et al., 2025a; Woźniak et al., 2024; Zhu et al., 2024; Kong et al., 2025), where a shared set of lightweight parameters (e.g., adapters or conditional batch normalization modules) is optimized to generalize across diverse user preferences. Conversely, the one-PEFT-per-user paradigm prioritizes individualized tuning through isolated parameter sets, enhancing personalization while preserving privacy by avoiding cross-user data leakage (Zhong et al., 2021; Peng et al., 2024; Tan et al., 2025; 2024). Beyond PEFT, RL-based approaches have gained traction for aligning models with user preferences through reward-driven optimization. Recent work has further formalized personalization as a multi-objective reinforcement learning (MORL) problem: methods like MORLHF (Wu et al., 2023) and MODPO (Zhou et al., 2024) train distinct reward models for different objectives and merge them during optimization, while alternative frameworks such as Personalized Soups (Jang et al., 2023), Reward Soups (Ramé et al., 2023), MOD (Shi et al., 2024), and PAD (Chen et al., 2025) dynamically combine policies from multiple trained models during the decoding phase. This series of work requires significant computational resources and relies on static datasets, which cannot meet the real-time personalization needs of users.
@@
Methods like Linear Alignment (Gao et al., 2024)and Contrastive Decoding (Li et al., 2023) pioneer logit steering mechanisms that dynamically adjust token distributions during generation. By computing differential signals between personalized and generic outputs on edge devices, these approaches eliminate dependency on retraining while preserving privacy through localized computation. CoS (He et al., 2025) follows this approach by using an adaptable parameter $\lambda$ to scale the impact of personal information, thereby achieving controllable personalization. Amulet (Zhang et al., 2025b) uses online learning algorithms to iteratively optimize the logits distribution, precisely aligning with the user’s preferences. However, this series of methods requires explicitly transmitting user information to the LLM, which poses a risk of privacy leakage (Yan et al., 2025) and hinders the practical deployment of these methods.
@@
The computational constraints of edge devices have driven innovative paradigms in collaborative generation between cloud-based LLMs and local small language models. Existing collaborative generation approaches predominantly focus on two technical routes: (1) assistive alignment, where local SLMs are trained to augment the LLM’s user preference alignment during inference, and (2) reward-guided decoding, where lightweight reward models provide preference signals through reward-guided decoding. For instance, Aligner (Ji et al., 2024) leverages natural language feedback generated by the SLM to directly inject user-specific linguistic patterns into the LLM’s decoding process, while Expo (Zheng et al., 2025) achieves preference alignment through linear interpolation of layer-wise weights between LLMs and SLMs. Recent advancements further demonstrate that training specialized small reward models can effectively steer LLM outputs toward personalized objectives (Snell et al., 2024; Liu et al., 2025b). Among these, Proxy-tuning (Liu et al., 2024) and Cogensis (Zhang et al., 2024b) are closest to our design. Proxy-tuning modifies the LLM’s logits distribution by contrasting pre-trained and fine-tuned SLM outputs, whereas Cogensis employs a learned fusion network to combine logits from both models. In contrast to these existing methods, our proposed CoSteer framework establishes a completely tuning-free collaboration mechanism, eliminating the need for training.
To clearly illustrate the technical positioning of our work, we have created Table 3 to highlight the key differences between our approach and related studies.
@@ 3.1.2 Online learning
Inspired by the recent work Amulet (Zhang et al., 2025b), our study employs the widely studied Follow-The-Regularized-Leader (FTRL) (McMahan, 2011) as the online learning algorithm. The FTRL is the FTL algorithm with a regularizer term added, which significantly improves both algorithmic stability and convergence properties. Typically, the policy optimization process of FTRL at iteration $t$ can be mathematically represented as:
@@ 3.2 Decoding-time personalization via local delta steering
Series of recent works, including Contrastive Decoding (Li et al., 2023) and Linear Alignment (Gao et al., 2024), have demonstrated that the difference in logits before and after incorporating specific context into the model can serve as a direction for optimization (He et al., 2025; Zhang et al., 2025b). Thus from the online learning perspective, their utility function can be expressed as follows:
@@
where $s$ denotes the current token sequence, $P_{LLM}$$\pi$ denotes the LLM’s distribution, and $P_{SLM}$$\pi^{*}$ the SLM’s distribution. Hyperparameters $\alpha,\beta$ adjust the influence of each optimization component.
@@
By substituting Equation 4 into 5, and then 5 into 6, we can iteratively optimize the strategy. However, the current iterative computation incurs significant overhead, which is not feasible for our personalization scenarios that require rapid response. Therefore, we deduce the closed-form solution of Equation 6, as shown in Equation 7, to significantly reduce the additional computational cost. The derivation of the closed-form solution can be found in Appendix A.3.
@@ 3.3 CoSteer framework
4 Experiment
4.1 Tasks and datasets
To demonstrate the versatility of our CoSteer framework, we conduct extensive experiments on eight datasets spanning two major categories of personalization tasks: personalized content generation and preference alignment.
Personalized content generation
We utilize two established benchmarks Cogenesis (Zhang et al., 2024b) and Longlamp (Kumar et al., 2024). Cogenesis provides summarized user profiles and past experiences, with the aim of generating highly personalized content. We evaluate on its official test set. Longlamp focuses on personalized long-form generation, where each writing query includes ground truth responses and the user’s previous writing records. We augment each instance with the top-5 relevant historical records retrieved by bge-reranker-v2-m3 (Chen et al., 2024) as its personal context. We conduct experiments on the official test sets of three constituent tasks: abstract generation (Tang et al., 2008), review writing (Ni et al., 2019), and topic writing (Völske et al., 2017).
Preference alignment
We evaluate on four datasets: Helpsteer (Wang et al., 2023b), Truthful QA (Lin et al., 2022), UltraChat (Ding et al., 2023), and Personal Preference Eval (Gao et al., 2024). These datasets require generating content aligned with explicitly stated user preferences. Due to the large scale of these datasets, we randomly sample 200 test instances from each dataset.
4.2 Evaluation metrics
To ensure fairness in comparison, we employ task-specific metrics proposed by these datasets themselves for evaluation. For Cogenesis, we use GPT-4o-2024-08-06 to evaluate the response’s overall and personalized scores, averaging the results over five runs with temperature set to 0 to mitigate potential instability. For Longlamp, we employ ROUGE (Lin, 2004) and METEOR (Banerjee & Lavie, 2005) scores to measure content overlap and semantic alignment, supplementing them with human evaluation (detailed in Appendix B.6) to account for the potential limitations of these automatic metrics. For preference alignment tasks, following prior work (Zhang et al., 2025b; Zhong et al., 2024), we set user preferences to be concise, creative, uplifting, and verbose, and utilize the reward model ArmoRM-8B (Wang et al., 2024) to evaluate the degree to which the generations align with these preferences.
4.3 Settings and baselines
To verify the robustness of our framework, we evaluate five model pairs of varying scales and architectures, including (1) Qwen2.5-7B-Instruct with Qwen2.5-1.5B-Instruct (2) Qwen2.5-32B-Instruct with Qwen2.5-7B-Instruct (Team, 2024) (3) Llama3.1-8B-Instruct with Llama-3.2-1B-Instruct (et al., 2024) (4) Qwen3-0.6B with Qwen3-8B and (5) Qwen3-0.6B with Qwen3-32B (Yang et al., 2025). Detailed parameters and settings are presented in Appendix B.1. We establish the following critical baselines per pair:
  • SLM w/: SLMs augmented with personal context.
  • LLM w/o: Cloud-based LLMs without access to personal context.
Qwen 7B-1.5B
SLM w/o 6.63 6.21 36.48 17.74 27.57 20.40 10.39 10.39 25.21 11.09 17.51 .5441 .5154 .6104 .5632
SLM w/ 7.81 7.63 39.75 22.03 27.76 23.08 12.40 12.94 22.89 11.46 17.12 .6214 .6998 .7102 .6594
LLM w/o 8.00 7.63 39.81 20.53 25.56 30.15 14.04 17.71 27.64 11.93 21.49 .7058 .6931 .7039 .7183
CoSteer 8.44 8.50 42.98 23.61 28.20 32.72 15.92 20.36 25.93 12.38 22.84 .7915 .7608 .7404 .7885
LLM w/ 8.62 8.60 44.50 24.63 31.15 33.83 15.55 22.42 28.82 13.77 23.44 .8884 .8076 .7693 .8414
Qwen 32B-7B
SLM w/o 8.00 7.63 39.81 20.53 25.56 30.15 14.04 17.71 27.64 11.93 21.49 .7058 .6931 .7039 .7183
SLM w/ 8.62 8.60 44.50 24.63 31.15 33.83 15.55 22.42 28.82 13.77 23.44 .8884 .8076 .7693 .8414
LLM w/o 8.12 7.87 40.66 21.02 26.61 32.21 14.44 19.61 28.82 12.20 21.16 .7020 .6873 .7225 .7146
CoSteer 8.78 8.64 45.41 26.04 33.52 34.88 15.89 26.51 30.10 14.52 24.20 .8589 .8532 .7274 .8579
LLM w/ 8.83 8.76 43.33 23.47 30.10 34.65 15.74 22.77 30.73 14.20 24.25 .9017 .8193 .7912 .8538
Llama 8B-1B
SLM w/o 7.04 6.55 33.20 18.20 28.55 31.75 14.92 19.78 20.81 10.21 17.30 .6535 .6355 .5900 .6646
SLM w/ 7.69 7.52 39.81 21.53 30.11 32.36 15.02 22.06 20.17 10.58 18.64 .7981 .7908 .7037 .8065
LLM w/o 7.69 7.13 39.33 20.69 29.41 34.58 15.32 22.10 26.93 12.35 21.82 .7038 .6878 .6765 .7116
CoSteer 7.29 7.73 41.28 24.97 31.19 31.68 13.68 24.57 26.11 12.00 23.69 .8911 .8582 .7812 .8721
LLM w/ 8.61 8.44 43.91 23.93 32.01 36.39 15.95 23.56 30.54 14.02 23.81 .8869 .8298 .7909 .8551
4.4 Main results
Table 1 presents our core experimental findings. For clarity, here we present results from first three of our five model pair configurations and report the average performance across the four preference alignment tasks. The complete results, including for the remaining two model pairs (Table 6) and the detailed per-dataset preference alignment scores (Table 5), can be found in the Appendix B.2. Furthermore, we conducted paired t-tests to verify that our improvements are statistically significant, with a detailed analysis available in Appendix B.3.
Overall Performance
We conducted extensive experiments across two major task categories (personalized content generation and preference alignment), different model architectures (Llama and Qwen), and various model sizes (from 0.6B to 32B, see Table 1 and  6 ) . In the vast majority of settings, our method significantly outperforms key baselines: (1) cloud-based LLMs responding without access to personal context, and (2) local SLMs with or without such context. Remarkably, our performance often approaches or even exceeds the “near-upper-bound” achieved by LLMs with full context access. These results show that CoSteer enables cloud-based LLMs to generate personalized content using only locally stored user context. We further analyze the impact of task complexity and model scale in Appendix B.4, confirming the robustness and adaptability of our approach.
4.5 Comparison with alternative methods
While CoSteer is unique in its technical positioning as shown in Table 3, to comprehensively demonstrate its superiority, we further compare it against other alignment methods, particularly those that rely solely on localized SLMs. Among the various existing techniques, we report on a selection of methods that can be applied to Personalized Content Generation tasks and that outperform the SLM w/ and SLM w/o baselines. These methods include:
  1. 1.
    Linear Alignment (Gao et al., 2024) and Context Steering(He et al., 2025) : These two methods are formally equivalent to scaling the logit difference of the SLM (with and without personal context) and applying it to its own inference process. This constitutes an inference-time optimization performed directly on the SLM.
  2. 2.
    Supervised Fine-Tuning (SFT): This involves directly fine-tuning the localized SLMs using personal data. Although we argue that fine-tuning a separate model for each user and task is impractical in real-world scenarios due to privacy and data constraints, we report its performance here for a thorough academic comparison.
  3. 3.
    Proxy-tuning Liu et al. (2024): This method leverages the difference in the SLM’s logit distributions before and after the SFT process described above. This difference is directly added to the LLM’s logits to steer its output distribution.
We conduct experiments on Longlamp using the Qwen7B-1.5B configuration. Implementation details and parameter settings for these baselines are provided in the Appendix B.8. As shown in Table 2, the results again confirm the superiority of the CoSteer framework. Our method adeptly combines the performance advantage of localized SLMs in handling private data with the powerful general capabilities of cloud-based LLMs, achieving the best performance among all compared methods without requiring any training.
Abstract Review Writing
R-1 R-L MET R-1 R-L MET R-1 R-L MET
LA/CS 41.49 25.57 29.96 26.60 13.68 17.59 24.09 12.32 19.94
SFT 40.71 21.85 27.00 31.55 12.91 20.23 25.53 10.80 19.35
Proxy tuning 40.76 21.37 25.95 29.54 14.16 16.79 27.13 12.11 21.38
[dashed] Costeer 42.98 23.61 28.20 32.72 15.92 20.36 25.93 12.38 22.84
w/o alpha 41.10 21.78 26.09 30.51 14.40 17.47 27.41 12.32 21.95
w/o KL 42.52 22.76 27.04 31.44 14.43 18.70 26.47 13.12 22.82
Table 2: Performance comparison of our Costeer framework against other methods on LongLamp benchmark, with an ablation study on its key components. Among four methods, best results are marked in bold while the second-best are underlined.
4.6 Ablation Study and Parameter Analysis
Ablation study
We conduct an ablation study to validate the contributions of the key components in our framework. The SLM delta steering signal (the $\beta$ term in Eq. (4)) is the central mechanism of CoSteer. Our study therefore focuses on the impact of individually removing two other components: the LLM Alignment Term (the $\alpha$ term) and the KL Divergence Regularizer from the FTRL algorithm. The results in Table 2 proved their effectiveness. A broader ablation on the entire iterative FTRL (we name it LightCosteer) is discussed in Section 5.3.
Hyperparameters
We examine the impact of different hyperparameters on our method. We conduct experiments on abstract generation using the Qwen7B-1.5B configuration, and evaluate with the ROUGE-L metric. When studying one parameter, we keep all others at their default values. Results and analysis presented in Appendix B.9 and Figure 2 show consistently strong performance across a broad range of values, confirming that the default hyperparameter settings (provided in Appendix B.1) are both robust and well-calibrated.
5 Discussion: Practical Deployment Considerations
In Section 4 , we showed the empirical superiority of the CoSteer framework through comprehensive experiments and comparisons. In this section, we shift our focus to its practical viability, analyzing its feasibility and performance from various angles that simulate real-world conditions and constraints.
5.1 Robustness to Noisy User Context
In practical on-device applications, the user-specific context is often imperfect and susceptible to noise. To evaluate CoSteer’s stability in such scenarios, we conducted rigorous stress tests under two distinct noise conditions: (1) realistic noise, simulated by replacing the strong bge-m3 retriever with a weaker BM25 retriever, and (2) adversarial noise, where we intentionally provided completely irrelevant context from a disjoint task.
Our experiments, detailed in Appendix B.10, yield a crucial insight: CoSteer demonstrates remarkable resilience to noisy context. Under realistic noise, it consistently outperforms baselines and maintains its performance edge (Table 9). In the more extreme adversarial setting, the framework degrades gracefully rather than collapsing, indicating that LLM guidance prevents the framework from being completely misled by faulty context (Table 10). This robustness is vital for real-world deployment.
5.2 Generalization across Model Scales and Architectures
A key advantage of our framework is its ability to generalize across diverse model configurations, a critical feature for real-world deployment. We validate this generalization capability in two dimensions: model scale and architecture.
Generalization Across Model Scales
Our primary experiments (see Table 6) already demonstrate remarkable scale generalization by pairing a powerful 32B-parameter LLM with a highly compact 0.6B SLM (Qwen3-0.6B), one of the smallest state-of-the-art models available. Despite being over 50× smaller, this SLM effectively steers the LLM across all our testbeds, strongly substantiates the framework’s robustness to vast differences in model scale.
Collaboration Across Model Architectures
While pairing models from the same family is common, a practical framework should accommodate scenarios where the LLM and SLM differ in architecture and tokenizer. To this end, we investigated two strategies. The first, CoSteer_map, aligns models via vocabulary intersection, while effective, its applicability is limited when tokenizers are highly divergent. Therefore, we propose CoSteer_byte which operates on a shared byte representation, making it vocabulary-agnostic (Hayase et al., 2025). As shown in Table 11, both strategies successfully enable the SLM to guide the LLM, confirming that CoSteer is not confined to homogeneous model families. This flexibility demonstrates the framework’s generalization across architectures. Full implementation details are provided in Appendix 11.
5.3 Efficiency and Practical Variants
Finally, we analyze the efficiency and overhead of our framework. At a high level, a key advantage of CoSteer is its computational efficiency. Compared to methods that require concatenating long personal context to the LLM’s input (i.e., LLM w/), our framework offers significant computational savings. By processing the private context locally with an SLM, CoSteer substantially reduces the remote LLM’s workload. Our FLOPs estimation (Narayanan et al., 2021) in Table 14 shows that with a 1000-token context, the privacy-violating LLM w/ approach is approximately 3.3x more computationally expensive than CoSteer.
Although our closed-form solution is efficient, the full CoSteer framework still employs iterative optimization (e.g., $T=20$), which can introduce latency. Furthermore, each generation step requires communication between the local device and the server. To address these practical overheads, we propose and evaluate two streamlined variants:
  • LightCoSteer: To minimize computational overhead, this variant eliminates the iterative process by setting the number of optimization iterations $T=1$. This simplifies the framework to a single-step logit adjustment, providing a lightweight yet effective alternative.
  • AdaCoSteer: To reduce communication overhead, this variant adaptively deactivates steering. We observe that as generation proceeds, SLM and LLM predictions often converge. AdaCoSteer leverages this by terminating the steering process once the LLM’s token confidence exceeds a set threshold for a few consecutive steps (Song et al., 2025), allowing the LLM to complete the generation on its own.
The implementation details and a full analysis of the results are presented in Appendix B.12. Our findings show that these variants create a valuable performance-efficiency spectrum, allowing users to select the optimal configuration that balances effectiveness and resource constraints for diverse application requirements.
6 Conclusion
In this work, we introduced CoSteer, a collaborative framework that enables real-time, privacy-preserving personalization for LLMs via localized delta steering. Our approach leverages an on-device SLM to guide a cloud-based LLM, effectively incorporating user context without transmitting sensitive data or requiring fine-tuning. Extensive experiments demonstrate that CoSteer offers a practical and effective solution for harmonizing the powerful general capabilities of LLMs with user-specific context and privacy.
Appendices Contents
Appendix A Methodology Detail
A.1 Related work
As shown in 3, CoSteer is unique as it is both collaborative and completely tuning-free. By formulating the steering as a local online learning problem, it allows for a lightweight, highly adaptive, and privacy protection framework.
Table 3: Our Technical position. The main body of the table uses “Yes/No” for clarity. Explanations for specific “No” entries are as follows: aRequires a trained reward model. bRequires a fine-tuned smaller language model (SLM). cRequires a trained fusion network.
Method
Training Free? Weak-to-Strong Collaborative? No Additional Modules?
Linear Alignment (Gao et al., 2024) Yes No Yes
Contrastive Decoding (Li et al., 2023)
Context Steering (He et al., 2025)
Amulet (Zhang et al., 2025b)
PAD (Chen et al., 2025) Noa No No
Drift (Kim et al., 2025)
Proxy-tuning (Liu et al., 2024) Nob Yes Yes
Cogenesis (Zhang et al., 2024b) Noc Yes No
CoSteer (Ours) Yes Yes Yes
A.2 Framework
Our CoSteer framework workflow is shown in Algorithm 1